OpenStudy (anonymous):

The curve whose equation is y=ax^3 + bx^2 + cx + d has a point of inflexion at (-1,4), a turning point when x=2, and it passes through the point (3,-7). Find the values of a, b, c and d.

OpenStudy (anonymous):

a=1 b=-3 c=-6 d=11 i think this is the answer.

OpenStudy (anonymous):

would you please explain it?

OpenStudy (anonymous):

u need to differentiate y with respect to x. point of inflexion and turning point hav dy/dx=0 then u can manipulate the infos given to get b,c,d in term of a. so, u let a=1, u wil get the others.

OpenStudy (anonymous):

I think a=3/11 b=9/11 c=-72/11 d=-23/11

OpenStudy (anonymous):

the point of inflection implis the 2nd derivative is 0 not the first

OpenStudy (anonymous):

really? i thought because the tangent is zero, so it also implies the first derivative is zero. idk, i got it wrong. lol.. thanks twhite.

OpenStudy (anonymous):

hmmm...got it twhitelsu....thanks both of you... :)